OpenStudy (anonymous):

Which of the points C(−6, 3) or D(3, 0) is closer to the point E(−2, 1): 1)Point C(−6, 3) is closer to point E(−2, 1). or 2)Point D(3, 0) is closer to point E(−2, 1). or 3) Points C(−6, 3) and D(3, 0) are the same distance from point E(−2, 1)

OpenStudy (anonymous):

distance between points: \[d(E->C)=\sqrt{(-2-(-6))^{2}+(1-3)^{2}}=\sqrt{16+4}=\sqrt{20}\]

OpenStudy (anonymous):

\[d(D->E)=\sqrt{(3-(-3))^{2}+(0-1)^{2}}=\sqrt{25+1}=\sqrt{26}\]

OpenStudy (anonymous):

so, point C is closer to point E